# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ Abstract.
We initiate the study of magnitude theory of real hyperplane arrangements. Magnitude is a cardinality-like invariant of metric spaces or enriched categories measuring the effective size. Its categorification, the magnitude homology, is a more powerful invariant. For a real hyperplane arrangement, or more generally, an oriented matroid, the tope graph encapsulates considerable amount of information. Since tope graphs are equipped with the shortest path metric, we feed them to the magnitude and magnitude homology machinery to derive new invariants of real hyperplane arrangements. We prove some structural results of the magnitude of arrangements, including reciprocity, palindromic numerator and denominator. For magnitude homology of arrangements, we give combinatorial descriptions in small length and prove that tope graphs are diagonal if and only if the arrangement is Boolean. We present a face decomposition of magnitude homology, using which we obtain a combinatorial formula of diagonal magnitude Betti numbers. Many open problems are posted for future study. In particular, we conjecture that magnitude and magnitude homology of arrangements are determined by the intersection lattice.
@@ 1. Introduction
Let $\mathcal{A}$ be an arrangement of hyperplanes in $\mathbb{R}^{\ell}$ all passing through the origin. The hyperplanes in $\mathcal{A}$ cut $\mathbb{R}^{\ell}$ into a disjoint union of chambers. Let $\mathrm{Ch}(\mathcal{A})$ denote the set of chambers. For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ as the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$.
The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$, and two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane. The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$. Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, isometric to a subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ equipped with the Hamming distance. For an oriented matroid $\mathcal{M}$, the tope graph $\mathcal{T}(\mathcal{M})$ can be defined similarly.
Although seemingly elementary, the tope graphs turn out to be quite informative.
@@
We refer to Figure 1 of [32] for a detailed diagram of implications. In particular, if two arrangements have isomorphic tope graphs, then they have homoemorphic complexified complements, isomorphic intersection lattices and isomorphic filtered Varchenko–Gelfand algebras. Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
@@
In this paper, we feed tope graphs into the emerging machinery of magnitude theory. Originally introduced by Leinster for enriched categories and metric spaces [24, 21], magnitude acts as a cardinality-like invariant measuring the “effective size” of a space. For finite graphs [25], it is a rational function defined via the inverse of the $q$-distance matrix, which is intimately related to the classical Varchenko matrix in the context of arrangements. The categorification of this concept, magnitude homology, developed by Hepworth and Willerton for graphs [16], and by Leinster and Shulman for enriched categories and metric spaces [23], provides a powerful bigraded homology theory that detects profound metric features.
We initiate the study of magnitude and magnitude homology specifically tailored to the metric geometry of real hyperplane arrangements. Our motivation is two-fold. The first is that we introduce a new rational function invariant and a new bigraded homology invariant of real arrangements. As a special class of graphs, tope graphs of arrangements (or oriented matroids) possess many unique features, such as antipodal symmetry and organized face decomposition, that enable us to contribute to the theory of magnitude with new beautiful examples.
At the decategorified level, we demonstrate that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, exhibits remarkable structural results. We prove a reciprocity formula relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, and establish that its reduced rational form consists of palindromic numerator and denominator whose poles are roots of unity. Supported by extensive computational evidence generated via SageMath on the Grünbaum catalogue [14], we further propose a conjectural Face Decomposition formula, which serves as a discrete $q$-analogue to the classical decomposition of the Euler characteristic of a zonotope.
At the homological level, by exploiting the metric structure of gated subgraphs within partial cubes, which naturally arise from the faces of the arrangement, we systematically dissect the magnitude homology of $\mathcal{T}(\mathcal{A})$. After providing explicit combinatorial descriptions of the homology groups for small lengths, we establish a strict topological dichotomy for arrangements. We prove that the tope graph is diagonal, meaning its off-diagonal magnitude homology strictly vanishes, if and only if the arrangement is Boolean. We also prove a face decomposition of magnitude homology of arrangements, and as an application, we derive a formula of diagonal magnitude Betti numbers in terms of the intersection lattice.
@@
The paper is organized as follows. Section 2 collects necessary recollections from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition. Section 3 investigates the basic properties and structural rigidity of the magnitude rational function. Section 4 explores magnitude homology, presenting the Boolean diagonality theorem and face decomposition theorem. Finally, Section 5 posts several open problems and conjectures for future study. Results of many computer experiments with SageMath form the appendix.
Convention: When we talk about a topological property/construction of a poset, we mean that of its order complex.
@@ 2.1. Metric graph theory
In this paper, all graphs we mention are supposed to be finite simple undirected graphs unless otherwise specified. For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from vertex $x\in V$ to vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$ and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$. We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ is the minimal length of walks from $x$ to $y$. If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$. Then $(G,d)$ is a generalized metric space. In particular, when $G$ is connected, that is there is a walk between any pair of vertices, $(G,d)$ is a metric space. We call the maximal value of $d$ the diameter of $G$, denoted by $\mathrm{diam}(G)$.
A subgraph $H$ of $G$ can be equipped with a metric in two ways. One is to restrict $d_{G}$ to $H$ and the other is $d_{H}$ on $H$ as a graph itself. We say $H$ is an isometric subgraph of $G$ if the two metrics coincide $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk. Given $x,y\in V$, let us define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$. It is a poset equipped with partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ traversing $u$ before traversing $v$. Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$. A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$. More generally, vertices $u_{1},\ldots,u_{k}$ form a strict linearly ordered chain in $x,y]_{G}$, that is $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$ if and only if
$d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ excluding $x$ and $y$.
We say a subgraph $H$ of $G$ is convex in $G$ if for any $x,y\in V(H)$, any minimal walk from $x$ to $y$ traverses only vertices in $H$, equivalently $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$. Apparently, any convex subgraph is an isometric subgraph, but the converse is false.
@@
Let $G$ be a connected graph, $H$ an isometric subgraph of $G$ and $x\in V(G)$. We say $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that for any $v\in V(H)$,
@@
Such $u$ must be unique since the above equation says $u$ is the closest vertex to $x$ in $H$. We may call $u$ the gate of $H$ with respect to $x$. Intuitively, any minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate. We say $H$ is a gated subgraph of $G$ is it is gated with respect to all vertices of $G$. In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection. Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
Proposition 2.2 (Proposition A.1 of [2]).
@@
Suppose $H$ is gated in $G$. Choose any minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$ and let $z$ be an arbitrary vertex in this walk, we need to show $z\in V(H)$. Since $H$ is gated, we have
@@
In the scope of this paper, the homomorphisms between graphs we shall use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, which are maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$. Note that gate projections are distance non-increasing.
@@ 2.2. Graph magnitude
See Leinster [25] for details. For a finite simple graph $G=(V,E)$, define the Zeta matrix $Z_{G}=Z_{G}(q)=[q^{d(x,y)}]_{x,y\in V}$, where $q^{\infty}=0$ by convention. Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$. The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ then can be expressed as a power series in $\mathbb{Z}[\![q]\!]$. In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
@@
In general, the sum of all entries of a matrix $A$ is equal to $\mathbf{1}^{T}A\mathbf{1}$, where $\mathbf{1}$ is the column vector of appropriate size with all entries equal to $1$. If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$. For graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$, in other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$. Then the magnitude
@@
There is a simple case that we can compute magnitude easily.
Proposition 2.3 (Lemma 3.2 of [25]).
@@
The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and thus denoted by $s=s(g)$. Then it is easily seen that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$. We conclude $\mathrm{Mag}(G)=\#V/s$. ∎
@@
Let $K_{n}$ be the complete graph on $n$ vertices and $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Then
@@
Magnitude behaves as cardinality in the following sense.
Proposition 2.5 (Lemmas 3.5 and 3.6 of [25]).
Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ the disjoint union and cartesian product of $G$ and $H$ respectively. Then
$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),~\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
Under some conditions, magnitude satisfies inclusion-exclusion.
Theorem 2.6 (Theorem 4.9 of [25]).

### 01 May 2026 v3

@@ Abstract.
We define and study the magnitude and magnitude homology of a real hyperplane arrangement by regarding its tope graph as a metric space. We prove several structural results for the magnitude of arrangements, including a symmetry formula, palindromicity of the numerator and denominator, a face decomposition formula, and results on the sign pattern of the magnitude power series. For the magnitude homology of arrangements, we obtain combinatorial formulas for small lengths and show that it detects Boolean arrangements. We also lift the face decomposition formula to a homological decomposition and derive explicit formulas for the diagonal magnitude Betti numbers. Another notable feature is that the magnitude Euler characteristic satisfies a reciprocity theorem analogous to Ehrhart–Macdonald reciprocity. We conclude by presenting several conjectures. In particular, we conjecture that the magnitude homology of an arrangement is torsion-free and is determined by the intersection lattice.
@@ 1. Introduction
Let $\mathcal{A}$ be an arrangement of hyperplanes passing through the origin in $\mathbb{R}^{d}$. The hyperplanes of $\mathcal{A}$ divide $\mathbb{R}^{d}$ into regions called chambers. We denote the set of chambers by $\mathrm{Ch}(\mathcal{A})$. The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ and with edges joining two chambers separated by a single hyperplane. More generally, the tope graph is defined for an oriented matroid. Although its definition is elementary, the tope graph contains essentially all combinatorial information about a hyperplane arrangement:
@@
We refer to Figure 1 of [34] for a detailed diagram of implications. In particular, if two arrangements have isomorphic tope graphs, then they have homeomorphic complexified complements, isomorphic intersection lattices, and isomorphic filtered Varchenko–Gelfand algebras. Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
@@
In this paper, by applying the emerging theory of magnitude to tope graphs, we define the magnitude and magnitude homology of arrangements and initiate the study of these invariants. Magnitude is a cardinality-like invariant of enriched categories and metric spaces, defined by Leinster [24, 21]; for metric spaces, it can be interpreted as measuring their “effective size.” The magnitude of a graph is defined by regarding the graph as a metric space via the shortest-path metric [25]. The magnitude of a graph is a rational function defined using the inverse of the $q$-distance matrix; in the case of hyperplane arrangements, this matrix coincides with the classical Varchenko matrix. The notion of magnitude homology, a categorification of magnitude, was introduced for graphs by Hepworth and Willerton [17], and for enriched categories and metric spaces by Leinster and Shulman [23]. It is a bigraded homology theory that reflects deep metric properties.
Our motivation is twofold. First, magnitude and magnitude homology give new invariants of real hyperplane arrangements. What information about an arrangement is reflected by these new invariants remains mysterious and is worth exploring. Second, the tope graphs of arrangements form a class of graphs with many distinctive features, such as antipodal symmetry and decompositions into faces, and may therefore serve as a source of interesting examples in the magnitude theory of graphs.
At the decategorified level, we show that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, satisfies remarkable structural properties. We prove a symmetry relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, which reflects the antipodal symmetry of the tope graph. The reduced rational form of $\mathrm{Mag}(\mathcal{A},q)$ has palindromic numerator and denominator, and all its poles are roots of unity. Moreover, if we define the interior magnitude by $\mathrm{Mag}^{\circ}(\mathcal{A},q)=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q)$, then the following face decomposition formula holds:
$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
Here $\mathcal{F}(\mathcal{A})$ denotes the set of faces (covectors) of $\mathcal{A}$, and $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subset H\}$. This serves as a discrete $q$-analogue of the classical decomposition of the Euler characteristic of a zonotope. We then investigate the sign pattern of the magnitude power series. In particular, we show that for arrangements of rank at most $3$, the signs of the magnitude coefficients eventually alternate. We also present an example with non-alternating magnitude.
The magnitude homology $MH_{k,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is a more refined invariant. By analyzing the structure of gated subgraphs arising from the faces of an arrangement, we establish the basic properties of this invariant. First, we describe all magnitude homology groups of length $\ell\leq 2$ in terms of the combinatorics of the arrangement. Next, we show that magnitude homology detects Boolean arrangements: an arrangement is Boolean if and only if its magnitude homology is concentrated on the diagonal. Furthermore, we define the interior magnitude homology $MH^{\circ}_{k,\ell}(\mathcal{A})$, which is a categorification of the interior magnitude $\mathrm{Mag}^{\circ}(\mathcal{A})$, and lift the face decomposition formula for magnitude to the homological level. In particular, we obtain the Euler characteristic reciprocity
$\chi^{\circ}_{\ell+\#\mathcal{A}}(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}}\chi_{\ell}(\mathcal{A}),$
where $\chi_{\ell}(\mathcal{A})$ (resp. $\chi^{\circ}_{\ell}(\mathcal{A})$) is the Euler characteristic of $MH_{*,\ell}(\mathcal{A})$ (resp. $MH^{\circ}_{*,\ell}(\mathcal{A})$). This can be regarded as a new combinatorial reciprocity theorem analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31]. As an application of the homological face decomposition, we give a complete description of the diagonal magnitude Betti numbers in terms of the intersection lattice.
@@
The paper is organized as follows. Section 2 recalls the necessary material from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition. Section 3 investigates the basic properties and structure of the magnitude rational function. Section 4 focuses on the sign pattern of magnitude power series. Section 5 explores magnitude homology, presenting the Boolean diagonality theorem and the face decomposition theorem. Finally, Section 6 poses several open problems and conjectures for future study. The appendix contains results from numerous SageMath computer experiments.
Use of AI
In writing this paper, we used AI in the following ways.
  • A substantial part of the proof strategy for Theorem 3.7 was generated by ChatGPT-5.4 Pro.
  • Part of the proof strategy for Theorem 4.5 was generated by ChatGPT-5.5 Pro.
  • We used ChatGPT-5.5 Pro to proofread the manuscript for minor errors.
We emphasize, however, that we have not used any AI-generated text without careful verification. The text of this paper has been written under the full responsibility of the authors.
Convention
In this paper, all graphs considered are finite, simple, and undirected unless otherwise specified. When we discuss a topological property or construction of a poset, we mean that of its order complex. We shall use the term “arrangement” to mean a real central hyperplane arrangement.
@@ 2.1. Metric graph theory
For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from a vertex $x\in V$ to a vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$, and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$. We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ to be the minimum length among walks from $x$ to $y$. If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$. Then $(G,d)$ is a generalized metric space. In particular, when $G$ is connected, that is, when there is a walk between any pair of vertices, $(G,d)$ is a metric space. We call the maximum value of $d$ the diameter of $G$, denoted by $\operatorname{diam}(G)$.
A subgraph $H$ of $G$ can be equipped with a metric in two ways. One is the restriction of $d_{G}$ to $H$, and the other is the graph metric $d_{H}$ on $H$ itself. We say that $H$ is an isometric subgraph of $G$ if the two metrics coincide, that is, if $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk. For $x,y\in V$, define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$. We equip it with the partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ that traverses $u$ before $v$. Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$. A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$. More generally, vertices $u_{1},\ldots,u_{k}$ form a strictly linearly ordered chain in $x,y]_{G}$, that is, $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$, if and only if
(2.1) $d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ obtained by excluding $x$ and $y$.
We say that a subgraph $H$ of $G$ is convex in $G$ if, for any $x,y\in V(H)$, every minimal walk from $x$ to $y$ traverses only vertices in $H$; equivalently, $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$. Every convex subgraph is an isometric subgraph, but the converse is false.
@@
Let $G$ be a connected graph, let $H$ be an isometric subgraph of $G$, and let $x\in V(G)$. We say that $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that, for any $v\in V(H)$,
@@
Such a vertex $u$ must be unique, since the above equation says that $u$ is the closest vertex to $x$ in $H$. We call $u$ the gate of $H$ with respect to $x$. Intuitively, every minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate. We say that $H$ is a gated subgraph of $G$ if it is gated with respect to every vertex of $G$. In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection. Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
Proposition 2.2 ([2, Proposition A.1]).
@@
Suppose $H$ is gated in $G$. Choose a minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$, and let $z$ be an arbitrary vertex in this walk. We need to show that $z\in V(H)$. Since $H$ is gated, we have
@@
In this paper, the graph homomorphisms we use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, that is, maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$. Note that gate projections are distance non-increasing.
@@ 2.2. Graph magnitude
See Leinster [25] for details. For a finite simple graph $G=(V,E)$, define the Zeta matrix $Z_{G}=Z_{G}(q)=[q^{d(x,y)}]_{x,y\in V}$, where $q^{\infty}=0$ by convention. Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$. The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ can then be expressed as a power series in $\mathbb{Z}[\![q]\!]$. In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
@@
In general, the sum of all entries of a matrix $A$ is equal to $\mathbf{1}^{T}A\mathbf{1}$, where $\mathbf{1}$ is the column vector of appropriate size with all entries equal to $1$. If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$. For a graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$. In other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$. Then the magnitude is given by
@@
The magnitude of a vertex-transitive graph can be computed explicitly.
Proposition 2.3 ([25, Lemma 3.2]).
@@
The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and we denote it by $s=s(g)$. It is then easy to see that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$. We conclude $\mathrm{Mag}(G)=\#V/s$. ∎
@@
Let $K_{n}$ be the complete graph on $n$ vertices and let $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Their magnitudes are given by
@@
Magnitude behaves like cardinality in the following sense.
Proposition 2.5 ([25, Lemmas 3.5 and 3.6]).
Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ for the disjoint union and Cartesian product of $G$ and $H$, respectively. Then
$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),\quad\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
Under certain conditions, magnitude satisfies inclusion-exclusion.
Theorem 2.6 ([25, Theorem 4.9]).